0.9 gram aluminium reacts with dilute HCl to give H2 the volume of H2 evolved at STP is

Grade:12

1 Answers

ANIKET RAJ
5 years ago
equation
Al+HCl--AlCl3+H2
moles of Al=0.9/27=0.033
moles of h2=0.033
moles=vol given/22.4
vol=22.4*0.033=0.7392 litres=739.2ml